### v1

@@ 2. Armagh Observatory and Planetarium, College Hill, Armagh BT61 9DG, Northern Ireland, UK
Polarimetric and spectropolarimetric observations with FoReRo2: Instrument overview and standard stars monitoring††thanks: Based on data collected with the 2-m RCC telescope at Rozhen National Astronomical Observatory.
@@
Standard stars with a high degree of polarisation are an essential part of the polarimetric observations. In our sample, HD 204827 shows short-term variability in position angle, making it unsuitable for calibration. HD 183143 displays intrinsic polarisation variability but remains stable in position angle. For the first time, we present a statistical study of K and of the Serkowski law, including their mean value, standard deviation, and distribution.
In order to show the FoReRo2 polarimetric capabilities, we also present several examples. The recurrent nova RS Oph shows variability of Serkowski’s K parameter due to dust formation within the first few days after the 2021 outburst. The examples also include imaging polarimetry of comet C/2019 Y4 (ATLAS) and spectropolarimetric observations of the symbiotic star Z And.
@@
Polarisation is a key characteristic of electromagnetic waves, along with intensity and colour, which opens a new horizon of object characterisation.
@@
The interest in polarimetric and spectropolarimetric observations has increased in recent years, as reflected in several instrumental papers describing various instruments (e.g., WALOP-North (1.3-m telescope, Skinakas Observatory, 2024JATIS..10d4005K), DIPol-UF (Nordic Optical Telescope, 2021AJ....161...20P), CasPol (2.1-5m Jorge Sahade Telescope, Complejo Astronómico El Leoncito, Argentina, 2019JATIS...5b8002S), DIPOL-1 (Sierra Nevada Observatory, 90-cm telescope, 2024AJ....167..137O), SPARC4 (1.6-m telescope, Observatório do Pico dos Dias, 2024BASBr..35...44R), SPHERE/ZIMPOL (VLT, 2018A&A...619A...9S) and ToPol (1.04-m Omicron (West) C2PU facility, Calern Observatory, France) (ToPol)) now replaced by a copy of the DIPol-UF instrument used at NOT (2021AJ....161...20P). This trend is further evidenced by the growing number of observations of standard stars, which play a key role in polarimetric calibration (e.g. 2023A&A...677A.144B; 2024MNRAS.535.1586C; 2024A&A...684A.132M; 2020A&A...635A..46P; 2020BAAA...61B.183L). At higher energies, this renewed emphasis is exemplified by the launch of the Imaging X-ray Polarimetry Explorer (IXPE), which has reopened the X-ray polarimetry window (2016SPIE.9905E..17W).
@@
In this paper, we characterise the polarimetric performances of the 2-Channel-Focal-Reducer Rozhen (hereafter FoReRo2 or FR2 (2000KFNTS...3...13J)). This instrument is attached to the Ritchey–Chrétien (RC) focus of the 2-m telescope of the Bulgarian National Astronomical Observatory (BNAO) – Rozhen. We use a homogeneous dataset of observations of both zero and high polarisation standard stars and present some examples of scientific applications. We assess the stability and the accuracy of the instrument over nearly a decade of observations, and to provide reliable calibration data that can support future scientific use of FR2.
@@
In section 2 we review the FoReRo2 and its observing modes. In section 3, we present the observing strategy and the data reduction steps. Section 4 contains the results of observations of stars with high degrees of polarisation. Section 5 showcases a statistical study of and of polarised stars, spectropolarimetric observations of the recurrent nova RS Oph and the symbiotic star Z And, as well as polarimetric observations (imaging) of the comet C/2019 Y4 (ATLAS). These observations serve as examples to demonstrate FoReRo2’s operational flexibility and its capability to deliver high-quality polarimetric and spectropolarimetric data for a diverse range of astronomical targets.
2 The Two-Channel-Focal-Reducer Rozhen
@@
FoReRo2 was delivered to BNAO based on a contract between the Max-Planck Institute for Solar System Research and the Institute of Astronomy (IA) in 2004. Since that time, the FR2 has been used at the f/8 RC focus of the 2m telescope, where it reduces the original focal ratio to f/2.8, corresponding to an effective focal length of 5.6 m. The FR2 mounted at the telescope is shown in Fig. 1.
The FR2 is a multimode instrument offering a wide variety of observing capabilities:
broadband and narrowband imaging
long-slit, low-dispersion (R) spectroscopy
Fabry-Perot interferometry
imaging polarimetry
low-dispersion spectropolarimetry
@@
A detailed description of the instrument is presented in 2000KFNTS...3...13J. Since that time, several changes and upgrades of the FR2 have been made. Now only two lenses are available, one in the red and one in the blue channel, colour-corrected in the range of 587–1014 nm and 365–436 nm, respectively. A set of Sloan filters and an H filter were added to the set of filters listed in 2000KFNTS...3...13J.
@@
One important improvement of the polarimetric mode is a new Wollaston prism which uses the full aperture of the telescope now. Thus, the previously used quadruple imaging polarimeter, described in 1996Ap&SS.239..259G, was replaced with a dual-beam instrument for polarisation measurements. The Wollaston prism is mounted in a special holder, which has the capability of linear translation (see Fig. 2). This construction allows a quick switch between the different observing modes. In the case of polarimetric and spectropolarimetric observations, the Wollaston prism is inserted into the parallel light beam, and it is taken out of the beam when switching to other observing modes.
@@
The Wollaston Quartz prism is placed in the parallel beam, before the colour divider, thus it feeds both channels simultaneously. The split angle of guarantees separation without overlapping of the ordinary (o) and extraordinary (e) stripes, imaged on the blue and red CCD (see the top panel in Fig. 5).
@@
For several years after the delivery of the FR2 in BNAO, polarimetric measurements were possible in imaging mode only. In 2011, we modified the instrument in such a way that it can also work in spectropolarimetric mode. This was a turning point which opened the way for low-dispersion spectropolarimetry, and the first observation obtained in the spectropolarimetric mode was of comet C/2009 P1 (Garrad)111(http://nao-rozhen.org/observations/fr8.html).
@@
For several years, the polarimetric mode was used by rotating the whole FR2 to different position angles in order to obtain the necessary and sufficient set of observations needed for calculating the Stokes parameters Q and U. In February 2014, an ”Agreement for collaborative research actions between Armagh Observatory (AO) and Institute of Astronomy and National Astronomical Observatory (IA and NAO) at the Bulgarian Academy of Sciences (BAS)” was signed. In the framework of this collaboration, AO delivered to BNAO a /2 retarder plate (half-wave plate (HWP)) for joint polarimetric observations with the 2-m telescope. The design and characteristics of waveplates can be found in 2004JQSRT..88..319S. The waveplate used with FR2 is a Super-Achromatic True Zero-Order Waveplate 5 - APSAW-5. More details are presented on the webpage of the vendor222http://astropribor.com/super-achromatic-quarter-and-half-waveplate/. A dedicated, precise mechanical construction and electronic control system for the rotation of the retarder was designed and manufactured in the workshop of the Institute of Astronomy and BNAO. All components are mounted on a plate, which is shown in Fig. 3. On the bottom side of the plate (not seen in the figure), a gear is mounted that holds the rotation cage (the housing of the HWP has a clear aperture of 40 mm). A worm drive gear, rotated by a stepping motor, is coupled to the cage.
The diameter of the /2 retarder is 40 mm, and it is mounted immediately below the RC focal plane (marked as HWP in Fig 2). A better solution would be to mount the /2 retarder in the parallel beam, but as the diameter of the latter is 50 mm, such a solution would cause light losses greater than 35%.
@@
The rotation of the /2 retarder is controlled remotely with a positional accuracy better than . This accuracy was achieved with precise mechanical construction and control electronics consisting of an Arduino board fitted with a Stepper Motor Driver.
The Arduino board runs the motor control and precise positioning firmware, which is accessed over a serial interface by a Raspberry Pi single-board computer. The Raspberry Pi computer runs the user software to control the angle of the /2 retarder plate and is accessed remotely over the Local Area Network.
@@
In order to avoid any backslashes in the gears, the firmware is designed in such a way that it always approaches the needed position from one direction only. The absolute angle rotation is achieved using a precise optical null position indicator, which is turned on only during the rotation in order to avoid any light spill during the exposure. The null position indicator is also used for self-calibration.
@@
The delivered HWP had no mark showing the orientation of its fast axis. Therefore, we performed a laboratory experiment with a twofold aim: (1) to identify the orientation of the fast axis, and (2) to check the mechanical uniformity over a full cycle of rotation. For this experiment, a white screen, inclined at with respect to the optical path, was inserted above the RC focal plane. It was illuminated with a tungsten lamp, and the reflected light passed through the HWP. After that, the o- and e-beams were separated by the Wollaston prism. The contrast between the o- and e-beams was increased by putting a Glan prism in the RC focal plane, i.e. between the reflecting screen and the HWP. The Glan prism was rotated manually until the best contrast was achieved (at angle of the HWP). After that, the HWP was rotated in steps of , covering the full range of angles from to . The signal of the o- and e- beams, and , respectively, was measured in rectangles of equivalent size 5661 px, positioned around the centres of their respective image stripes. The result of this laboratory experiment is shown in Fig. 4. The three panels contain dependencies of the rotation angle of the fast axis of the HWP. In the top panel of this figure, the measured signals of both beams, and , are shown, as squares and triangles, respectively. The middle panel shows the calculated values of the quantity and its approximation with a function. The bottom panel shows the difference between the measured values and the fit. These residuals exhibit substantial fluctuations in uniformity during the full cycle of rotation. One relatively uniform region, close to the difference (data-fit) of , is seen in the range from to . We should note that this laboratory experiment is used with the FR2 only, not with the entire system telescope plus focal reducer.
@@
In 2018, both channels of the FR2 were equipped with new detectors. These are Andor CCD cameras iKon-L 936333https://andor.oxinst.com/products/ikon-ccd-cameras-for-long-exposure, with the new dual antireflection deep depletion ’BEX2-DD’ chip. One great advantage of the new cameras is the thermo-electric cooling, and another one, especially for the red channel, is the suppression of the fringing in the red and near-IR spectral region. The cameras comprise 20482048 quadratic pixels, each having a size of 13.5 m with a resulting scale of 0.497 arcsec/px and a FoV of about 17x17 arcmin (in imaging mode).
@@
For spectropolarimetric observations, we use a grism with 300 lines in the red channel. Two slit widths are available: 110 m and 200 m. Using the 110 m slit yields a resolving power of at , with a reciprocal spectral dispersion of 3 pixel-1. Over the past eight years (after recoating the 2 m mirror) in order to obtain more accurate results, we used only slitless spectropolarimetry (See Section 3.3.1 for details). In the slitless mode, the resolving power depends on the seeing.
@@
Raw images of the polarimetric (top: comet C/2019 Y4 (ATLAS)) and spectropolarimetric (bottom: Be/X-ray binary star X Per) mode of observations are presented in Fig. 5. For the polarimetric mode of observations, a mask of size 35 mm 4 mm is used. For slitless spectropolarimetric observation, a mask of size 4 mm 3 mm is used. Two spectra of Be/X-ray binary star X Per are presented in Fig. 5, where the emission line is visible. The spectra correspond to the o- and e- beam, respectively.
@@ 3.1 Observations and basic reduction steps
We obtain polarised spectra and polarised images at eight HWP angles apart.
Standard stars with zero degrees of polarisation were used to correct the instrumental polarisation. Standard stars with a high degree of polarisation were used to correct the position angle. Polarisation standards were observed with the same instrumental setup as astronomical objects on the same nights.
The spectra are reduced using IRAF (1993ASPC...52..173T) in the standard way, including bias removal and wavelength calibration. For the spectropolarimetric data reduction, our research group created dedicated IRAF scripts. For better comparison with catalogue values in the V filter, we used the extracted and from spectra using the IRAF SBAND procedure.
The process of spectropolarimetric data reduction includes the following: bias subtraction; extracting 1D spectra; wavelength calibration; beam swapping technique; corrections for instrumental polarisation, chromatism of the HWP, and position angle correction. The beam swapping technique (BST) (2009PASP..121..993B) is the core of the sequence of reduction steps. We use it to minimise the instrumental polarisation.
We tested spectropolarimetric data reduction with and without the use of flat fields. Our tests show that in both cases, with and without flat fields, the results obtained are indistinguishable within the precision of our data. This result corresponds to the result obtained by other authors (2019JATIS...5b8002S).
3.2 Beam swapping technique
A beam-swapping technique is used for polarimetric data processing. The values of F and , where and F are the fluxes of the o- and e- beams, respectively, are obtained for the different HWP angles. We define , where corresponds to the HWP angle.
The following formulae are used to calculate the Stokes parameters (2009PASP..121..993B):

### v2

@@ 2. Armagh Observatory and Planetarium, College Hill, Armagh BT61 9DG, Northern Ireland, UK
Polarimetric and spectropolarimetric observations with FoReRo2: Instrument overview and standard star monitoring††thanks: Based on data collected with the 2-m RCC telescope at Rozhen National Astronomical Observatory.
@@
Standard stars with a high degree of polarisation are an essential part of the polarimetric observations. In our sample, HD 204827 shows short-term variability in the position angle (PA), making it unsuitable for calibration. HD 183143 displays intrinsic polarisation variability, but remains stable in terms of the PA. For the first time, we present a statistical study of the K and parameters of Serkowski’s law, including the mean value, standard deviation, and distribution.
To demonstrate FoReRo2’s polarimetric capabilities, we present several examples. The recurrent nova RS Oph shows the variability of Serkowski’s K parameter, which is due to dust formation within the first few days after the 2021 outburst. These examples also include the imaging polarimetry of comet C/2019 Y4 (ATLAS) and spectropolarimetric observations of the symbiotic star Z And.
@@
Polarisation is a key characteristic of electromagnetic waves, together with intensity and colour, opening up a new horizon in the field of object characterisation. The interest in polarimetric and spectropolarimetric observations has increased in recent years, as reflected in several instrumental papers describing various instruments (e.g., WALOP-North, 1.3-m telescope, Skinakas Observatory, Kypriotakis et al. 2024; DIPol-UF, Nordic Optical Telescope, Piirola et al. 2021; CasPol, 2.15-m Jorge Sahade Telescope, Complejo Astronómico El Leoncito, Argentina, Sosa et al. 2019; DIPOL-1, Sierra Nevada Observatory, 90-cm telescope, Otero-Santos et al. 2024; SPARC4, 1.6-m telescope, Observatório do Pico dos Dias, Rodrigues et al. 2024; SPHERE/ZIMPOL, VLT, Schmid et al. 2018; and ToPol, 1.04-m Omicron (West) C2PU facility, Calern Observatory, France Devogèle et al. 2017). We note that the latter has been replaced by a copy of the DIPol-UF instrument used at NOT (Piirola et al., 2021). This trend is further evidenced by the growing number of observations of standard stars, which play a key role in polarimetric calibration (e.g. Blinov et al. (2023); Cotton et al. (2024); Mandarakas et al. (2024); Piirola et al. (2020); López et al. (2020)). At higher energies, this renewed emphasis is exemplified by the launch of the Imaging X-ray Polarimetry Explorer (IXPE), which has reopened the X-ray polarimetry window (Weisskopf et al., 2016).
@@
In this paper, we characterise the polarimetric performance of the 2-Channel-Focal-Reducer Rozhen (hereafter, FoReRo2 or FR2 (Jockers et al., 2000)). This instrument is attached to the Ritchey–Chrétien (RC) focus of the 2-m telescope of the Bulgarian National Astronomical Observatory (BNAO) – Rozhen. We use a homogeneous dataset of observations of both zero- and high-polarisation standard stars and present a set of examples of scientific applications in this context. We assess the stability and the accuracy of the instrument over nearly a decade of observations, and provide reliable calibration data that can support future scientific use of FR2.
@@
In Section 2, we review the FoReRo2 and its observing modes. In Section 3, we present the observing strategy and the data reduction steps. Section 4 contains the results of observations of stars with high degrees of polarisation. Section 5 showcases a statistical study of and of polarised stars, spectropolarimetric observations of the recurrent nova RS Oph and the symbiotic star Z And, as well as polarimetric observations (imaging) of the comet C/2019 Y4 (ATLAS). These observations serve as examples to demonstrate FoReRo2’s operational flexibility and its capability to deliver high-quality polarimetric and spectropolarimetric data for a diverse range of astronomical targets.
2 Two-Channel-Focal-Reducer Rozhen
@@
FoReRo2 was delivered to BNAO based on a contract between the Max-Planck Institute for Solar System Research and the Institute of Astronomy (IA) in 2004. Since that time, the FR2 has been used at the f/8 RC focus of the 2m telescope, where it reduces the original focal ratio to f/2.8, corresponding to an effective focal length of 5.6 m. The FR2 mounted at the telescope is shown in Fig. 1. FR2 is a multi-mode instrument offering a wide variety of observing capabilities:
broad-band and narrow-band imaging;
long-slit, low-dispersion (R) spectroscopy;
Fabry-Perot interferometry;
imaging polarimetry;
low-dispersion spectropolarimetry.
@@
A detailed description of the instrument is presented in Jockers et al. (2000). Since that time, several changes and upgrades of the FR2 have been made. Now only two lenses are available, one in the red and one in the blue channel, colour-corrected in the range of 587–1014 nm and 365–436 nm, respectively. A set of Sloan filters and an H filter were added to the set of filters listed in Jockers et al..
@@
One important improvement of the polarimetric mode is a new Wollaston prism, which uses the full aperture of the telescope now. Thus, the previously used quadruple imaging polarimeter, described in Geyer et al. (1996), was replaced with a dual-beam instrument for polarisation measurements. The Wollaston prism is mounted in a special holder, which has the capability of linear translation (see Fig. 2). This construction allows a quick switch between the different observing modes. In the case of polarimetric and spectropolarimetric observations, the Wollaston prism is inserted into the parallel light beam, and it is taken out of the beam when switching to other observing modes.
@@
The Wollaston Quartz prism is placed in the parallel beam, before the colour divider; thus it feeds both channels simultaneously. The split angle of guarantees separation without overlapping of the ordinary (o) and extraordinary (e) stripes, imaged on the blue and red CCD (see the top panel in Fig. 5).
@@
For several years after the delivery of the FR2 in BNAO, polarimetric measurements were possible in imaging mode only. In 2011, we modified the instrument in such a way that it can also work in spectropolarimetric mode. This was a turning point which opened the way for low-dispersion spectropolarimetry and the first observation obtained in the spectropolarimetric mode was of comet C/2009 P1 (Garrad)111http://nao-rozhen.org/observations/fr8.html.
@@
For several years, the polarimetric mode was used by rotating the whole FR2 to different angles to obtain the necessary and sufficient set of observations needed for calculating the Stokes parameters Q and U. In February 2014, the Agreement for collaborative research actions between Armagh Observatory (AO) and Institute of Astronomy and National Astronomical Observatory (IA and NAO) at the Bulgarian Academy of Sciences (BAS) was signed. In the framework of this collaboration, AO delivered to BNAO a /2 retarder plate (half-wave plate (HWP)) for joint polarimetric observations with the 2-m telescope. The design and characteristics of waveplates can be found in Samoylov et al. (2004). The waveplate used with FR2 is a Super-Achromatic True Zero-Order Waveplate 5 - APSAW-5. More details are presented on the webpage of the vendor222http://astropribor.com/super-achromatic-quarter-and-half-waveplate/. A dedicated, precise mechanical construction and electronic control system for the rotation of the retarder was designed and manufactured in the workshop of the Institute of Astronomy and BNAO. All components are mounted on a plate, which is shown in Fig. 3. On the bottom side of the plate (not seen in the figure), a gear is mounted that holds the rotation cage (the housing of the HWP has a clear aperture of 40 mm). A worm drive gear, rotated by a stepping motor, is coupled to the cage.
@@
The diameter of the /2 retarder is 40 mm, and it is mounted immediately below the RC focal plane (marked as HWP in Fig 2). A better solution would be to mount the /2 retarder in the parallel beam, but as the diameter of the latter is 50 mm, such a solution would cause light losses greater than 35%. The rotation of the /2 retarder is controlled remotely with a positional accuracy better than . This accuracy was achieved with precise mechanical construction and control electronics consisting of an Arduino board fitted with a stepper motor driver.
@@
The Arduino board runs the motor control and precise positioning firmware, which is accessed over a serial interface by a Raspberry Pi single-board computer. The Raspberry Pi computer runs the user software to control the angle of the /2 retarder plate and is accessed remotely over the local area network. To avoid any backslashes in the gears, the firmware is designed in such a way that it always approaches the needed position from one direction only. The absolute angle rotation is achieved using a precise optical null position indicator, which is turned on only during the rotation in order to avoid any light spill during the exposure. The null position indicator is also used for self-calibration.
@@
The delivered HWP had no mark showing the orientation of its fast axis. Therefore, we performed a laboratory experiment with a twofold aim: (1) to identify the orientation of the fast axis; and (2) to check the mechanical uniformity over a full cycle of rotation. For this experiment, a white screen, inclined at with respect to the optical path, was inserted above the RC focal plane. It was illuminated with a tungsten lamp, and the reflected light passed through the HWP. Afterwards, the o- and e-beams were separated by the Wollaston prism. The contrast between the o- and e-beams was increased by putting a Glan prism in the RC focal plane (i.e. between the reflecting screen and the HWP). The Glan prism was rotated manually until the best contrast was achieved (at angle of the HWP). Afterwards, the HWP was rotated in steps of , covering the full range of angles from to . The signal of the o- and e- beams, and , respectively, was measured in rectangles of equivalent size 5661 px, positioned around the centres of their respective image stripes. The result of this laboratory experiment is shown in Fig. 4. The three panels contain dependencies of the rotation angle of the fast axis of the HWP. In the top panel of this figure, the measured signals of both beams, and , are shown, as squares and triangles, respectively. The middle panel shows the calculated values of the quantity and its approximation with a function. The bottom panel shows the difference between the measured values and the fit. These residuals exhibit substantial fluctuations in uniformity during the full cycle of rotation. One relatively uniform region, close to the difference (data-fit) of , is seen in the range from to . We should note that this laboratory experiment is used with the FR2 only, not with the entire system telescope plus focal reducer.
@@
In 2018, both channels of the FR2 were equipped with new detectors. These are Andor CCD cameras iKon-L 936333https://andor.oxinst.com/products/ikon-ccd-cameras-for-long-exposure, with the new dual antireflection deep depletion ’BEX2-DD’ chip. Two significant advantages of the new cameras are the thermo-electric cooling and suppression of the fringing in the red and near-IR spectral region, which is especially useful for the red channel. The cameras comprise 20482048 quadratic pixels, each having a size of 13.5 m with a resulting scale of 0.497 arcsec/px and a FoV of about 17x17 arcmin (in imaging mode).
@@
For spectropolarimetric observations, we use a grism with 300 lines in the red channel. Two slit widths are available: 110 m and 200 m. Using the 110 m slit yields a resolving power of at , with a reciprocal spectral dispersion of 3 pixel-1. Over the past eight years (after re-coating the 2 m mirror) to obtain more accurate results, we only used slitless spectropolarimetry (see Section 3.4 for details). In the slitless mode, the resolving power depends on the seeing.
@@
Raw images of the polarimetric (top: comet C/2019 Y4 (ATLAS)) and spectropolarimetric (bottom: Be/X-ray binary star X Per) mode of observations are presented in Fig. 5. For the polarimetric mode of observations, a mask of size 35 mm 4 mm is used. For slitless spectropolarimetric observations, a mask of size 4 mm 3 mm was used. Two spectra of the Be/X-ray binary star X Per are presented in Fig. 5, where the emission line is visible. The spectra correspond to the o- and e- beam, respectively.
@@ 3.1 Observations and basic reduction steps
We obtained polarised spectra and polarised images at eight HWP angles apart. Standard stars with zero degrees of polarisation were used to correct the instrumental polarisation. Standard stars with a high degree of polarisation were used to correct the PA. Polarisation standards were observed with the same instrumental setup as astronomical objects on the same nights.
The spectra are reduced using IRAF (Tody, 1993) in the standard way, including bias removal and wavelength calibration. For the spectropolarimetric data reduction, our research group created dedicated IRAF scripts. For better comparison with catalogue values in the V filter, we used the extracted and from spectra using the IRAF SBAND procedure.
The process of spectropolarimetric data reduction includes the following: bias subtraction, extracting 1D spectra, wavelength calibration, beam-swapping technique, and corrections for instrumental polarisation, chromatism of the HWP, and the PA. The beam-swapping technique (BST) (Bagnulo et al., 2009) was at the core of the sequence of reduction steps. We used it to minimise the instrumental polarisation. We tested spectropolarimetric data reduction with and without the use of flat fields. Our tests show that in both cases, with and without flat fields, the results obtained are indistinguishable within the precision of our data. This result is in agreement with the result obtained by other authors (Sosa et al., 2019).
3.2 Beam-swapping technique
A beam-swapping technique is used for polarimetric data processing. The values of F and , where and F are the fluxes of the o- and e- beams, respectively, obtained for the different HWP angles. We can define , where corresponds to the HWP angle. The following formulae are used to calculate the Stokes parameters (Bagnulo et al., 2009):